# Instrumental Variable Approach to Estimating Individual Causal Effects in N-of-1 Trials: Application to ISTOP Study

### v2

@@ Abstract
Keywords: “N-of-1” trial; instrumental variable; confounding; causal inference; time series; mHealth; wearables.
@@ 4.3.1 Models
where $\mathbf{f}(\cdot)$ and $\mathbf{g}(\cdot)$ denote any general functions including variable selection, and $h(\cdot)$ is a link function. $w_{tj}$ denotes the set of all possible measured confouders at period $t$ day $j$. We limit the coefficient $s$ in the second equation to be either 1 or $-1$, i.e., $s$ = {1,-1} and denotes the sign. In Appendix A.1, we show that $u_{tj}$ represents a summary of unmeasured confounders. For Gaussian models it is adequate to assume the same $u_{tj}$ for both equations and the magnitude of the coefficient is not important.
@@ A.1 Confounding Coefficient
Consider two scenarios, where in scenario (A) (Figure S1) we assume an unmeasured confounder $\tilde{u}_{tj}\sim N(0,\tilde{\sigma}_{u}^{2})$ and its effect on $X_{tj}^{\dagger}$ and $Y_{tj}^{\dagger}$ are $a$ and $b$ respectively; in scenario (B) (Figure S1) we assume an unmeasured confounder ${u}_{tj}\sim N(0,|ab|\tilde{\sigma}_{u}^{2})$ and its effect on $X_{tj}^{\dagger}$ and $Y_{tj}^{\dagger}$ are 1 and $sgn(ab)$. We argue that two scenarios are equivalent with regard to estimation models and $u_{tj}$ is considered a summary of the confounding effect. A mathematical explanation is provided below.
Scenario (A) is represented by the following structural models:
@@
If regressing $Y_{tj}^{\dagger}$ directly on $X_{tj}$, the bias of estimating $\beta$ due to their correlation is $\frac{ab\tilde{\sigma}_{u}^{2}}{\sigma_{x}^{2}}\text{E}(X_{tj}^{\dagger}|X_{tj})$, a quantity dependent on the product of the coefficients and its variance $ab\tilde{\sigma}_{u}^{2}$ given fixed $\sigma_{x}$ and $x_{tj}$. In other words, the individual effect $a$ or $b$ or the variance $\tilde{\sigma}_{u}^{2}$ do not individually affect or contribute to the bias. Therefore we can replace $a\tilde{u}_{tj}$ and $b\tilde{u}_{tj}$ with a new term $u_{tj}$ = $\sqrt{ab}\tilde{u}_{tj}$ to represent a summary of the confounding effect, i.e., scenario (B) represented by the following structural models:
$X_{tj}^{\dagger}$ $=\alpha_{0}+\alpha_{1}r_{t}+{u}_{tj}+\tilde{\epsilon}^{x}_{tj},$
$Y_{tj}^{\dagger}$ $=\beta_{0}+\beta_{1}X_{tj}+sgn(ab){u}_{tj}+\tilde{\epsilon}^{y}_{tj},$ (18)
where ${u}_{tj}\sim N(0,ab\tilde{\sigma}_{u}^{2}).$

### v3

@@ Abstract
Keywords: “N-of-1” trial; Bayesian; instrumental variable; confounding; causal inference; time series; mHealth; wearables.
@@ 4.3.1 Models
where $\mathbf{f}(\cdot)$ and $\mathbf{g}(\cdot)$ denote any general functions including variable selection, and $h(\cdot)$ is a link function. $w_{tj}$ denotes the set of all possible measured confouders at period $t$ day $j$. We limit the coefficient $s$ in the second equation to be either 1 or $-1$, i.e., $s$ = {1,-1} and denotes the sign of the product of the unknown effect of the confounders to $X_{tj}$ and $Y_{tj}$. In Appendix A.1, we show that $u_{tj}$ represents a summary of unmeasured confounders. For Gaussian models it is adequate to assume the same $u_{tj}$ for both equations and the magnitude of the coefficient is not important.
@@ 6 Discussion
7 Acknowledgments
The authors express our gratitude to Dr. Gregory M. Marcus and his team for providing access to the data in the I-STOP-AFib trials. I-STOP-AFib trials were funded in part through a Patient-Centered Outcomes Research Institute (PCORI) Award (PPRND-1507- 31321). Christopher H. Schmid and Tao Liu are partially supported by Institutional Development Award Number U54GM115677 from the National Institute of General Medical Sciences of the National Institutes of Health, which funds Advance Rhode Island Clinical and Translational Research (Advance RI-CTR). The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
8 Conflicts of Interest
All authors declare that they have no conflicts of interest.
@@ A.1 Confounding Coefficient
Consider two scenarios, where in scenario (A) (Figure S1) we assume an unmeasured confounder $\tilde{u}_{tj}\sim N(0,\tilde{\sigma}_{u}^{2})$ and its effect on $X_{tj}^{\dagger}$ and $Y_{tj}^{\dagger}$ are $a$ and $b$ respectively; in scenario (B) (Figure S1) we assume an unmeasured confounder ${u}_{tj}\sim N(0,|ab|\tilde{\sigma}_{u}^{2})$ and its effect on $X_{tj}^{\dagger}$ and $Y_{tj}^{\dagger}$ are 1 and $sgn(ab)$. We argue that the two scenarios cause the same magnitude of bias in estimating the causal effect parameter ($\beta_{1}$ in Eq.(17)). Given that the two scenarios cannot be distinguished without knowing $\tilde{u}_{tj}$ or $u_{tj}$, we propose to use $u_{tj}$ as a summary of the unconfounding effects in our simulation studies. Scenario (A) is represented by the following structural models:
@@
If regressing $Y_{tj}^{\dagger}$ directly on $X_{tj}$, the bias of estimating $\beta_{1}$ due to correlation between $Y_{tj}^{\dagger}$ and $X_{tj}^{\dagger}$ is $\frac{ab\tilde{\sigma}_{u}^{2}}{\sigma_{x}^{2}}\text{E}(X_{tj}^{\dagger}|X_{tj})$, a quantity dependent on the product of the coefficients $a$ and $b$, and the variance $\tilde{\sigma}_{u}^{2}$ given fixed $\sigma_{x}$ and $x_{tj}$. In other words, the effect $a$ or $b$ or the variance $\tilde{\sigma}_{u}^{2}$ do not affect the bias individually; instead, they contribute to it as a composite. Therefore we can replace $a\tilde{u}_{tj}$ and $b\tilde{u}_{tj}$ with a new term $u_{tj}$ = $\sqrt{|ab|}\tilde{u}_{tj}$ to represent a summary of the confounding effect, i.e., scenario (B) represented by the following structural models:
$X_{tj}^{\dagger}$ $=\alpha_{0}+\alpha_{1}r_{t}+{u}_{tj}+\tilde{e}^{x}_{tj},$
$Y_{tj}^{\dagger}$ $=\beta_{0}+\beta_{1}X_{tj}+sgn(ab){u}_{tj}+\tilde{e}^{y}_{tj},$ (18)
where ${u}_{tj}\sim N(0,ab\tilde{\sigma}_{u}^{2}).$ With Eq.(18), it follows that
$\text{E}(Y_{tj}^{\dagger}|X_{tj})$
$=\beta_{0}+\beta_{1}X_{tj}+\frac{ab\tilde{\sigma}_{u}^{2}}{\sigma_{x}^{2}}(\text{E}(X_{tj}^{\dagger}|X_{tj})-\alpha_{0}-\alpha_{1}r_{t}),$ (19)
yielding the same bias as in Eq.(17) of estimating $\beta_{1}$.